# What makes math problems hard for reinforcement learning: a case study

### v1

@@ 7. Language Modeling
In this section, we discuss a model for the “language” of balanced presentations. Each presentation with two relators is a sequence made of six letters, also known as “tokens” in the nomenclature of Natural Language Processing, i.e. $x$, $y$, $x^{-1}$, and $y^{-1}$, and two “stop tokens”: one that separates two relators of a presentation and another that marks the end of a presentation. Given this vocabulary $V$ of six tokens, we can ask what is the probability $p(t_{1},\dots,t_{N})$ for $t_{i}\in V$ of the occurrence of a specific presentation in the space of all balanced presentations. Using the chain rule of probability theory,
@@
Over the last few years, Transformer models have shown great success in modeling human-understandable languages to the extent that these models can create text almost indistinguishable from that of a human expert. Specifically, the architecture used for modeling language is the auto-regressive “decoder-only” Transformer, which we review in detail in Subsection 7.1. In Subsection 7.2, we discuss the method with which we generate the dataset required for training the model. Finally, in Subsection 7.3, we share details of some insights we learned from this process.
@@ 7.1. Transformers: a review
Given an input sequence $t_{1},t_{2},\dots,t_{N}$, a decoder-only transformer predicts the probability distribution $p(t\mid t_{1},t_{2},\dots,t_{N})$ over the set $V$ of tokens of size $n_{\text{vocab}}$. The probability is computed by applying the softmax function to the logits $T(t)$, which are estimated by applying the following sequence of operations.151515 The softmax function, $\text{Softmax}\colon\mathbb{R}^{n}\to(0,1)^{n}$, is defined as $\text{Softmax}(x)_{i}=e^{x_{i}}/\sum\limits_{j=1}^{n}e^{x_{j}}$. First, assign to each token in the vocabulary a distinct label in the range $1,2,\dots,n_{\text{vocab}}$; re-writing the original sequence as a sequence of integers. We will label these integers also as $t_{i}$. Next, write the sequence in terms of “one-hot encoded vectors”, i.e. a matrix $t\in\mathbb{R}^{N\times n_{\text{vocab}}}$ such that
@@
and embed the sequence in a $d_{\text{model}}$-dimensional vector space,161616 Here, $t$ and all $x_{j}$ are two-dimensional tensors. Hence, it is appropriate to apply tensors of linear transformations to them. Often in a transformer architecture, these operations are of the form $\mathbb{1}\otimes\cdots$; in these cases, we drop the identity transformation and simply write the operation as $\cdots$. For example, $\mathbb{1}\otimes W_{U}$, $\mathbb{1}\otimes W^{m}_{I}$, $\mathbb{1}\otimes W^{m}_{O}$, etc. In this case, we will sometimes write $W_{U}$, $W^{m}_{I}$, $W^{m}_{O}$ respectively, assuming it is clear from the context and the dimensionality of these matrices that they are tensored with identity transformations.
@@ 7.2. Training and Evaluation Datasets
We paid special attention to ensure that our dataset contains presentations of a wide range of lengths so as not to bias our model towards learning trends specific to any fixed length. To this end, we devised an algorithm (Algorithm 6 in Appendix C) that creates an almost uniform distribution over the lengths of the presentations. (See Figure 17.) We set aside $10\%$ of our entire data for validation.
@@
Figure 17. Percentage of presentations in various ranges of lengths. Percentages were computed independently for the two subsets of the dataset, corresponding to presentations that are AC-equivalent to GS-solved and GS-unsolved presentations. We used Algorithm 6 from Appendix C to ensure the almost-uniform distribution depicted here.
@@ 7.3. Results
A randomly initialized model with the initialization scheme given in [Radford2019LanguageMA] has a cross entropy loss of $-\ln(1/n_{\text{vocab}})\approx 1.7918$. With training, we could achieve a validation loss of $0.7337$.171717 We tuned the hyperparameters a little but it is quite likely that one can achieve a better performing model with more hyperparameter tuning. Similarly, more training data will necessarily help with the performance. We trained a Transformer model with hyperparameters given in Appendix A. We used the untrained and the trained model to get the embeddings of all $1190$ presentations of the Miller–Schupp series with $n\leq 7$ and $\operatorname{length}(w)\leq 7$. We used t-SNE to project these embedding vectors to a plane [JMLR:v9:vandermaaten08a]. The plots are shown in grid in Figure 18.
Each row of Figure 18 corresponds to a fixed value of $n$. The left (resp. right) column depicts t-SNE projections of embeddings obtained by an untrained (resp. trained) model. t-SNE dependence on a distance measure: it learns to map vectors that are closer together in the higher-dimensional space, with respect to this distance measure, close together in the plane [JMLR:v9:vandermaaten08a]. We used cosine simiarity between embedding vectors as the distance measure for our plots. We note that the GS-solved and GS-unsolved presentations seem to cluster much more in the plots in the right column. This indicates that a trained Transformer model is able to distinguish between GS-solved and GS-unsolved presentations to a good extent, albeit not perfectly.181818 Note also that t-SNE admits a hyperparameter known as “perplexity”, and the projections learned by t-SNE depend on the hyperparameter [wattenberg2016how]. Thus, in general, t-SNE plots must be interpreted with care. The plots shown in Figure 18 were all made with the perplexity value of $30$. We checked however that the clusters of GS-solved and GS-unsolved presentations continue to exists at a broad range of perplexity values.
@@
    There likely exists an invariant at the level of the “language“ of the balanced presentations that distinguishes GS-solved vs GS-unsolved presentations.

### v2

@@ 6. Language modeling
One of the main messages in this paper is that dynamically learning the notion of hardness is of paramount importance to the success of RL agents in problems with extremely long horizons and sparse rewards. One way of approaching this goal is to build a larger AI system in which RL model is accompanied by another component that is good at recognizing hard vs easy instances, AC presentations in our case study. Therefore, as a prerequisite, it is natural to explore data analysis tools that can do such identification. In this section, we explore how well Language Models can do this job, and in the next section we explore alternative data analysis tools based on the structure of the AC graph, whose vertices represent balanced presentations and edges represent basic AC moves.
The approach via language modeling appears very natural because balanced presentations are described by “words” in group generators, and this is something that language models learn remarkably well. Specifically, each presentation with two relators is a sequence made of six letters, also known as “tokens” in the nomenclature of Natural Language Processing, i.e. $x$, $y$, $x^{-1}$, and $y^{-1}$, and two “stop tokens”: one that separates two relators of a presentation and another that marks the end of a presentation. Given this vocabulary $V$ of six tokens, we can ask what is the probability $p(t_{1},\dots,t_{N})$ for $t_{i}\in V$ of the occurrence of a specific presentation in the space of all balanced presentations. Using the chain rule of probability theory,
@@
Over the last few years, Transformer models have shown great success in modeling human-understandable languages to the extent that these models can create text almost indistinguishable from that of a human expert. Specifically, the architecture used for modeling language is the auto-regressive “decoder-only” Transformer, which we review in detail in Section 6.1. In Section 6.2, we discuss the method with which we generate the dataset required for training the model. Finally, in Section 6.3, we share details of some insights we learned from this process.
@@ 6.1. Transformers: a review
Given an input sequence $t_{1},t_{2},\dots,t_{N}$, a decoder-only transformer predicts the probability distribution $p(t\mid t_{1},t_{2},\dots,t_{N})$ over the set $V$ of tokens of size $n_{\text{vocab}}$. The probability is computed by applying the softmax function to the logits $T(t)$, which are estimated by applying the following sequence of operations.202020 The softmax function, $\text{Softmax}\colon\mathbb{R}^{n}\to(0,1)^{n}$, is defined as $\text{Softmax}(x)_{i}=e^{x_{i}}/\sum\limits_{j=1}^{n}e^{x_{j}}$. First, assign to each token in the vocabulary a distinct label in the range $1,2,\dots,n_{\text{vocab}}$; re-writing the original sequence as a sequence of integers. We will label these integers also as $t_{i}$. Next, write the sequence in terms of “one-hot encoded vectors”, i.e. a matrix $t\in\mathbb{R}^{N\times n_{\text{vocab}}}$ such that
@@
and embed the sequence in a $d_{\text{model}}$-dimensional vector space,212121 Here, $t$ and all $x_{j}$ are two-dimensional tensors. Hence, it is appropriate to apply tensors of linear transformations to them. Often in a transformer architecture, these operations are of the form $\mathbb{1}\otimes\cdots$; in these cases, we drop the identity transformation and simply write the operation as $\cdots$. For example, $\mathbb{1}\otimes W_{U}$, $\mathbb{1}\otimes W^{m}_{I}$, $\mathbb{1}\otimes W^{m}_{O}$, etc. In this case, we will sometimes write $W_{U}$, $W^{m}_{I}$, $W^{m}_{O}$ respectively, assuming it is clear from the context and the dimensionality of these matrices that they are tensored with identity transformations.
@@ 6.2. Training and evaluation datasets
We paid special attention to ensure that our dataset contains presentations of a wide range of lengths so as not to bias our model towards learning trends specific to any fixed length. To this end, we devised an algorithm (Algorithm 6 in Appendix D) that creates an almost uniform distribution over the lengths of the presentations. (See Figure 16.) We set aside $10\%$ of our entire data for validation.
@@
Figure 16. Percentage of presentations in various ranges of lengths. Percentages were computed independently for the two subsets of the dataset, corresponding to presentations that are AC-equivalent to GS-solved and GS-unsolved presentations. We used Algorithm 6 from Appendix D to ensure the almost-uniform distribution depicted here.
@@ 6.3. Results
A randomly initialized model with the initialization scheme given in [Radford2019LanguageMA] has a cross entropy loss of $-\ln(1/n_{\text{vocab}})\approx 1.7918$. With training, we could achieve a validation loss of $0.7337$.222222 We tuned the hyperparameters a little but it is quite likely that one can achieve a better performing model with more hyperparameter tuning. Similarly, more training data will necessarily help with the performance. We trained a Transformer model with hyperparameters given in Appendix A. We used the untrained and the trained model to get the embeddings of all $1190$ presentations of the Miller–Schupp series with $n\leq 7$ and $\operatorname{length}(w)\leq 7$. We used t-SNE to project these embedding vectors to a plane [JMLR:v9:vandermaaten08a]. The plots are shown in grid in Figure 17.
Each row of Figure 17 corresponds to a fixed value of $n$. The left (resp. right) column depicts t-SNE projections of embeddings obtained by an untrained (resp. trained) model. t-SNE dependence on a distance measure: it learns to map vectors that are closer together in the higher-dimensional space, with respect to this distance measure, close together in the plane [JMLR:v9:vandermaaten08a]. We used cosine simiarity between embedding vectors as the distance measure for our plots. We note that the GS-solved and GS-unsolved presentations seem to cluster much more in the plots in the right column. This indicates that a trained Transformer model is able to distinguish between GS-solved and GS-unsolved presentations to a good extent, albeit not perfectly.232323 Note also that t-SNE admits a hyperparameter known as “perplexity”, and the projections learned by t-SNE depend on the hyperparameter [wattenberg2016how]. Thus, in general, t-SNE plots must be interpreted with care. The plots shown in Figure 17 were all made with the perplexity value of $30$. We checked however that the clusters of GS-solved and GS-unsolved presentations continue to exists at a broad range of perplexity values.
@@
    There likely exists an invariant at the level of the “language” of the balanced presentations that distinguishes GS-solved vs GS-unsolved presentations.